A park playground slide is d = 5.0m in length and is at an angle of Ø = 20° with respect to the ground. A boy of mass mb = 20 kg starts from rest at the top of the slide. The coefficient of sliding friction for the slide is µk = 0.2. How long does the boy take to slide down the ramp?       


1
Expert's answer
2021-05-08T14:37:37-0400
"a=g(\sin{20}-\mu \cos{20})\\d=0.5at^2\\d=0.5g(\sin{20}-\mu \cos{20})t^2
\\5=0.5(9.8)(\sin{20}-0.2 \cos{20})t^2\\t=2.6\ s"